# Adaptive Query Prompting for Multi-Domain Landmark Detection

### v1

@@ Abstract
Medical landmark detection is crucial in various medical imaging modalities and procedures. Although deep learning-based methods have achieve promising performance, they are mostly designed for specific anatomical regions or tasks. In this work, we propose a universal model for multi-domain landmark detection by leveraging transformer architecture and developing a prompting component, named as Adaptive Query Prompting (AQP). Instead of embedding additional modules in the backbone network, we design a separate module to generate prompts that can be effectively extended to any other transformer network. In our proposed AQP, prompts are learnable parameters maintained in a memory space called prompt pool. The central idea is to keep the backbone frozen and then optimize prompts to instruct the model inference process. Furthermore, we employ a lightweight decoder to decode landmarks from the extracted features, namely Light-MLD. Thanks to the lightweight nature of the decoder and AQP, we can handle multiple datasets by sharing the backbone encoder and then only perform partial parameter tuning without incurring much additional cost. It has the potential to be extended to more landmark detection tasks. We conduct experiments on three widely used X-ray datasets for different medical landmark detection tasks. Our proposed Light-MLD coupled with AQP achieves SOTA performance on many metrics even without the use of elaborate structural designs or complex frameworks.
@@ I INTRODUCTION
Landmark detection plays an important role in various of medical image analysis tasks [1, 2, 3]. It aims to localize anatomical keypoints in medical images such as X-ray images. Manual annotation is highly time-consuming, labor-intensive and typically requires a high level of expertise. Automatic annotation methods can significantly reduce the required labor and enable the analysis of large-scale datasets [4]. Despite the fact that deep learning-based methods have shown great promise on in the field of medical landmark detection, a significant limitation of many existing models is their task-specific nature. These models are typically tailored and refined to optimize performance within the confines of their designated tasks. Zhu et al. propose GU2Net [5] and DATR [6] as universal models for multi-domain landmark detection. However, these methods focus on elaborate structural designs to obtain better model performance, and require to train the entire model when handling new tasks, which is highly inefficient.
@@
To this end, we propose Adaptive Query Prompting (AQP), a novel prompting method for muti-domain learning. Figure 1 gives an overview of our method in contrast to typical task specific methods and existing universal models. The basic idea is to efficiently adapt a frozen large foundation model to many downstream tasks with the minimum extra trainable parameters [7]. We keep the pre-trained model untouched, and instead learn a set of prompts that dynamically instruct model to solve corresponding tasks, it leverages the representative features from pre-trained models. The prompts are selected via a query-based mechanism. Specifically, prompts and keys are paired in a shared memory space called prompt pool, then we design a query function to get proper prompts paired with the keys, which are selected based on the query function. Our proposed AQP can select the appropriate prompt for various input, and does not require complex design. The prompt pool, optimized jointly with the supervised loss, serves to ensure that shared prompts encode shared knowledge for effective knowledge transfer, while unshared prompts encode task-specific knowledge crucial for preserving model plasticity. The trainable parameters of our method is much fewer, thus being able to adapt to different tasks more efficiently.
In order to handle multiple landmark detection datasets without much extra cost, we employ a plain vision transformer as backbone to encode features and then adopt multiple lightweight decoders to deal with different tasks [8, 9, 10, 11]. For each iteration, we randomly sample instances from multiple training datasets and feed them into the backbone and the decoders to estimate the heatmaps corresponding to each dataset. Compared to previous methods, AQP enjoys the structure simplicity and can adapt well to different tasks through fine-tuning.
@@ II-A Vision Transformer for Medical Landmark Detection
Recently, vision transformers [12, 13] have shown great potential in many vision tasks [14, 15, 16]. For instance, Xiao et al. [17] propose a lightweight Transformer-embedded network for vertebrae landmark detection, which has yielded promising results in this task. PRTR [18] incorporates both transformer encoders and decoders to gradually refine the locations of the estimated keypoints in a cascade manner. However, Most of these approaches employ a convolutional neural network (CNN) as a backbone, followed by the integration of a transformer with intricate architectures to enhance the extracted features and capture the relationships among the keypoints. They either necessitate CNNs for feature extraction or demand meticulous design of transformer architectures. In the contrast, ViTPose [13] employs a plain vision transformer [12] as the backbone with a simple decoder and obtains SOTA performance on representative benchmarks without elaborate designs. Follow its success, we introduced this framework into medical landmark detection.
@@ II-B Prompt Tuning
Prompting was originally introduced in the field of natural language processing (NLP) [19]. Shyam et al. [20] illustrates robust generalization to downstream transfer learning tasks, even in few-shot or zero-shot settings, utilizing manually selected prompts within GPT-3. Recently, the concept of prompting [21, 22] has been extended to vision tasks. Sandler et al. [22] introduces memory tokens, which are sets of trainable embedding vectors for each transformer layer. These traditional prompt tuning methods design task-specific prompt functions to instruct pre-trained models perform corresponding tasks conditionally [19, 23], so that the language model gets additional information about the task. However, crafting an effective prompting function poses a challenge, which require heuristic strategies. In this work, instead of designing a prompting function, we learn a set of prompts stored in the prompt pool. A query function is used to dynamically select suitable prompts tailored to different inputs. All that remains is the design of an effective query function, which is much easier than design prompts for different tasks. We can utilize the entire pre-trained model as a frozen feature extractor to obtain the query features.
@@ III-A1 ViTPose
ViTPose [13] is a transformer-based network for human pose estimation. Similar to traditional CNN backbone [24], ViTPose captures multi-stale features via several stages. Differently, each stage is built via the feature embedding layers and vision transformer blocks [12, 25]. ViTPose adopt plain vision transformer with masked image modeling pre-training backbones, then pre-training the backbones using MAE [26] on MS COCO [27] dataset and a combination of MS COCO and AI Challenger [28] respectively by random masking 75% patches from the images and reconstructing those masked patches. In this work, we use the pre-trained weights to initialize the backbone of our model.
@@ III-A2 Lightweight decoder
Unlike the existing methods which use elaborate designs to decode the heatmaps from extracted features, we avoid introducing fancy but complex modules to keep the structure as simple as possible. We simply append several decoder layers after the transformer backbone to estimate the heatmaps. As shown in Figure 2, we directly utilize bilinear interpolation to upsample the feature maps by a factor of 4, followed by applying a Rectified Linear Unit (ReLU) and a convolutional layer with a kernel size of 3 × 3 to obtain the heatmaps, i.e.,
$K=Conv_{3\times 3}(Bilinear(ReLu(F_{out}))),$ (2)
@@ III-A3 Prompt-based learning
Given an input of 2D image $I\in\mathbb{R}^{H\times W\times C}$ and a pre-trained vision transformer (ViT) backbone $f=f_{r}\circ f_{e}$ (excluding the decoder head), where $f_{r}$ denotes transformer blocks [12], and $f_{e}$ represents the patch embedding layer. Images are reshape to a sequence of flattened 2D patches $I^{p}\in\mathbb{R}^{N\times(C\cdot S)}$, where $N$ is the number of patches. A image patch $I^{p}$ is projected to a D-dimension feature through the patch embedding layer, i.e., $f_{e}:\mathbb{R}^{N\times(C\cdot S)}\rightarrow\mathbb{R}^{S\times D}$ , where $D$ is the embedding dimension. The extracted features are denoted as $x_{e}$, and $x_{e}=f_{e}(I_{p})\in\mathbb{R}^{S\times D}$. When employing prompt tuning methods, prompts can be fed into model in various ways. For example, a prompt $P_{e}\in\mathbb{R}^{L_{p}\times D}$ is considered as an additional token concatenated to the embedding future $x_{e}$, where $L_{p}$ is the token length. In a remarkable work called Explicit Visual Prompting (EVP) [7], prompts are first adjusted by a tunable adapter, then added to the extracted embedding feature $x_{e}$ through element-wise addition. This kind of prompt tuning typically yields better results because it learns explicit prompts. In this work, we adopt a approach similar to EVP.
@@ III-B Overview
As illustrated in Figure 2, Light-MLD employs a transformer-based backbone as encoder, which captures multi-stale features via several vision transformer blocks, then followed by a series of lightweight decoders to decode the corresponding landmarks for different tasks. Figure 3 gives a overview our proposed Adaptive Query Prompting (AQP). Prompts are selected from the prompt pool to instruct the model during training process.
@@ III-C Adaptive query prompting
In our proposed adaptive query prompting, there are three main components, namely the Prompt pool, the Prompt query mechanism and the Adaptor.
@@ III-C1 Prompt pool
where $N$ represents total number of prompts. $P_{j}\in\mathbb{R}^{L_{p}\times D}$ is a single prompt with the same embedding size $D$ as $x_{e}$ and token length $L_{p}$. $x_{e}$ is extracted embedding feature projected by the embedding layer.
@@ III-C2 Prompt query mechanism
Let $\gamma:\mathbb{R}^{D_{k}}\times\mathbb{R}^{D_{k}}\rightarrow\mathbb{R}$ represent a function that scores the match between the query and prompt key (in this work we adopt cosine distance). When presented with an input $x$, we use $q(x)$ to retrieve the top-M keys by straightforwardly solving the objective function:
@@ III-C3 Adaptor
$P^{i}_{f}=MLP_{up}(GELU(MLP^{i}_{tune}(P))),$ (5)
where $MLP_{up}$ represents an up-projection layer shared across all Adaptors, aimed at aligning the dimensions of transformer feature. $GELU$ denotes the GELU activation function [29]. $MLP_{tune}$ is a linear layer responsible for generating distinct prompts in each Adaptor. $P^{i}_{f}$ is the output prompt that attaches to each transformer layer. This design allows for strong adaptability while generating diverse prompts. With the adaptor, we can generate suitable prompts for all tasks in a unified paradigm even when using different models or backbones.
@@ III-D2 Loss function
$\begin{split}\mathcal{L}=\frac{1}{2N}\sum_{j=0}^{N-1}||H^{j}-G^{j}||_{2}+\lambda\sum_{\mathbf{K}_{{\bm{x}}}}\gamma\left({q({\bm{x}}),{\bm{k}}_{s_{i}}}\right),\end{split}$ (6)
@@ IV-A1 Head
The head dataset is a widely-used public dataset for cephalometric landmark detection, which contains 400 X-ray images, provided in IEEE ISBI 2015 challenge [30]. It contains 19 landmarks for each X-ray image. For our experimentation, we designate the first 150 images for training and the remaining 250 images for testing purposes.
@@ IV-A2 Hand
The hand dataset comprises 909 X-ray images. We allocate the first 609 images for training and reserve the remaining 300 images for testing. Payer et al. [31] manually labeled a total of 37 landmarks.
@@ IV-A3 Chest
Models MRE Head SDR(%) MRE Hand SDR(%) MRE Chest SDR(%)
(mm) 2mm 2.5mm 3mm 4mm (mm) 2mm 4mm 10mm (px) 3px 6px 9px
Ibragimov et al. [32] 1.84 68.13 74.63 79.77 86.87 - - - - - - - -
Štern et al. [33] - - - - - 0.80 92.20 98.45 99.83 - - - -
Lindner et al. [34] 1.67 70.65 76.93 82.17 89.85 0.85 93.68 98.95 99.94 - - - -
Urschler et al. [35] - 70.21 76.95 82.08 89.01 0.80 92.19 98.46 99.95 - - - -
Payer et al. [36] - 73.33 78.76 83.24 89.75 0.66 94.99 99.27 $9$
U-Net [37]+ 12.45 52.08 60.04 66.54 73.68 6.14 81.16 92.46 93.76 5.64 51.67 82.33 90.67
GU2Net [5]+ 1.56 77.79 84.65 89.41 94.93 0.84 95.40 99.35 99.75 5.57 57.33 82.67 89.33
AQP (Ours)+ 1.54 76.39 84.43 89.61 95.02 0.76 96.27 99.45 99.79 5.61 56.78 83.46 90.04
Figure 4: Subjective results of the head and hand datasets. All images are randomly selected. The red points $\bullet$ are the landmarks predicted by our model while the green points $\bullet$ are the ground truth labels.
@@ IV-B Settings
[…]

### v2

@@ Abstract
Medical landmark detection is crucial in various medical imaging modalities and procedures. Although deep learning-based methods have achieve promising performance, they are mostly designed for specific anatomical regions or tasks. In this work, we propose a universal model for multi-domain landmark detection by leveraging transformer architecture and developing a prompting component, named as Adaptive Query Prompting (AQP). Transformer backbone architecture is suitable for our study due to its ability to capture long-range dependencies crucial in multi-domain landmark detection. Specifically, transformers excel at understanding global anatomical relationships, which span across entire images. Instead of embedding additional modules in the backbone network, we design a separate module to generate prompts that can be effectively extended to any other transformer network. In our proposed AQP, prompts are learnable parameters maintained in a memory space called prompt pool. The central idea is to keep the backbone frozen and then optimize prompts to instruct the model inference process. Furthermore, we employ a lightweight decoder to decode landmarks from the extracted features, namely Light-MLD. Thanks to the lightweight nature of the decoder and AQP, we can handle multiple datasets by sharing the backbone encoder and then only perform partial parameter tuning without incurring much additional cost. It has the potential to be extended to more landmark detection tasks. We conduct experiments on three widely used X-ray datasets for different medical landmark detection tasks. Our proposed Light-MLD coupled with AQP achieves SOTA performance on many metrics even without the use of elaborate structural designs or complex frameworks.
@@ I INTRODUCTION
Landmark detection plays an important role in various of medical image analysis tasks [Chen2024-dg, Gong2023GenerativeAF, Huang2023MRIS, chen11, liu2020fine, li1, li2]. It aims to localize anatomical keypoints in medical images such as X-ray images. Manual annotation is highly time-consuming, labor-intensive and typically requires a high level of expertise. Automatic annotation methods can significantly reduce the required labor and enable the analysis of large-scale datasets [medsam]. Despite the fact that deep learning-based methods have shown great promise on in the field of medical landmark detection [chen12, chen13, chen14, jiang2021deep, jiang2020geometry, zhang2022correction, li2020diagnosis, li2021hybrid, zhang1, li2023optimized, yang2024adaptive, huang2024deformmlp, li2022few, zhang2, zhang3, zhang4, li2022monocular, song2024local, li2023cee, zhang5, zhang6, zhang7, zhang8, zhang9, zhang10], a significant limitation of many existing models is their task-specific nature. These models are typically tailored and refined to optimize performance within the confines of their designated tasks. Zhu et al. propose GU2Net [yolo] and DATR [datr] as universal models for multi-domain landmark detection. However, these methods focus on elaborate structural designs to obtain better model performance, and require to train the entire model when handling new tasks, which is highly inefficient.
@@
To this end, we propose Adaptive Query Prompting (AQP), a novel prompting method for muti-domain learning. Figure 1 gives an overview of our method in contrast to typical task specific methods and existing universal models. The basic idea is to efficiently adapt a pre-trained large foundation model to many downstream tasks with the minimum extra trainable parameters [evp, jia2022visual, liu2023explicit2]. In the context of our multi-domain landmark detection task, AQP allows us to use the same transformer backbone for head, hand, and chest X-ray landmark detection by learning task-specific prompts. This is in contrast to conventional approaches that might require training separate models for each task or fine-tuning the entire model for each domain. We keep the pre-trained model untouched, and instead learn a set of prompts that dynamically instruct model to solve corresponding tasks, it leverages the representative features from pre-trained models. The prompts are selected via a query-based mechanism. Specifically, prompts and keys are paired in a shared memory space called prompt pool, then we design a query function to get proper prompts paired with the keys, which are selected based on the query function. Our proposed AQP can select the appropriate prompt for various input, and does not require complex design. The prompt pool, optimized jointly with the supervised loss, serves to ensure that shared prompts encode shared knowledge for effective knowledge transfer, while unshared prompts encode task-specific knowledge crucial for preserving model plasticity. The trainable parameters of our method is much fewer, thus being able to adapt to different tasks more efficiently.
In order to handle multiple landmark detection datasets without excessive extra training cost, we employ a plain vision transformer as backbone to encode features and then adopt multiple lightweight decoders to deal with different tasks [Chen2023ALF, Li2023HighResolutionDS, Luo2023DevignetHV, liu2023coordfill, chen15]. For each iteration, we randomly sample instances from multiple training datasets and feed them into the backbone and the decoders to estimate the heatmaps corresponding to each dataset. Compared to previous methods, AQP enjoys the structure simplicity and can adapt well to different tasks through fine-tuning.
@@ II-A Vision Transformer for Medical Landmark Detection
Recently, vision transformers [vit, vitpose] have shown great potential in many vision tasks [Chen2022ShadocnetLS, li2022wavenhancer, liu2024depth, chen10, chen7, liu2024dh]. For instance, Xiao et al. [bib3] propose a lightweight Transformer-embedded network for vertebrae landmark detection, which has yielded promising results in this task. PRTR [bib41] incorporates both transformer encoders and decoders to gradually refine the locations of the estimated keypoints in a cascade manner. However, Most of these approaches employ a convolutional neural network (CNN) as a backbone, followed by the integration of a transformer with intricate architectures to enhance the extracted features and capture the relationships among the keypoints. They either necessitate CNNs for feature extraction or demand meticulous design of transformer architectures. In the contrast, ViTPose [vitpose] employs a plain vision transformer [vit] as the backbone with a simple decoder and obtains SOTA performance on representative benchmarks without elaborate designs. Follow its success, we introduced this framework into medical landmark detection.
@@ II-B Prompt Tuning
Prompting was originally introduced in the field of natural language processing (NLP) [bib15]. Shyam et al. [bib6] illustrates robust generalization to downstream transfer learning tasks, even in few-shot or zero-shot settings, utilizing manually selected prompts within GPT-3. Recently, the concept of prompting [vpt, bib17] has been extended to vision tasks. Sandler et al. [bib17] introduces memory tokens, which are sets of trainable embedding vectors for each transformer layer. These traditional prompt tuning methods design task-specific prompt functions to instruct pre-trained models perform corresponding tasks conditionally [bib15, Chen2023MedPromptCP], so that the language model gets additional information about the task. However, crafting an effective prompting function poses a challenge, which require heuristic strategies. In this work, instead of designing a prompting function, we learn a set of prompts stored in the prompt pool. A query function is used to dynamically select suitable prompts tailored to different inputs. All that remains is the design of an effective query function, which is much easier than design prompts for different tasks. We can utilize the entire pre-trained model as a frozen feature extractor to obtain the query features.
@@ III-A1 ViTPose
ViTPose [vitpose] is a transformer-based network for human pose estimation. Similar to traditional CNN backbone [bib18], ViTPose captures multi-stage features via several stages. Differently, each stage is built via the feature embedding layers and vision transformer blocks [vit, attention]. ViTPose adopt plain vision transformer with masked image modeling pre-training backbones, then pre-training the backbones using MAE [mae] on MS COCO [coco] dataset and a combination of MS COCO and AI Challenger [bib22] respectively by random masking 75% patches from the images and reconstructing those masked patches. In this work, we use the pre-trained weights to initialize the backbone of our model.
@@ III-A2 Lightweight decoder
Unlike the existing methods which use elaborate designs to decode the heatmaps from extracted features, we avoid introducing fancy but complex modules to keep the structure as simple as possible. We simply append several decoder layers after the transformer backbone to estimate the heatmaps. As shown in Figure 2, we directly utilize bilinear interpolation to upsample the feature maps by a factor of 4, followed by applying a Rectified Linear Unit (ReLU) and a convolutional layer with a kernel size of $3\times 3$ to obtain the heatmaps, i.e.,
$K=\operatorname{Conv}_{3\times 3}(\operatorname{Bilinear}(\operatorname{ReLu}(F_{out}))),$ (2)
@@ III-A3 Prompt-based learning
Given an input of 2D image $I\in\mathbb{R}^{H\times W\times C}$ and a pre-trained vision transformer (ViT) backbone $f=f_{r}\circ f_{e}$ (excluding the decoder head), where $f_{r}$ denotes transformer blocks [vit], and $f_{e}$ represents the patch embedding layer. Images are reshape to a sequence of flattened 2D patches $I^{p}\in\mathbb{R}^{N\times(C\cdot S)}$, where $N$ is the number of patches. A image patch $I^{p}$ is projected to a D-dimension feature through the patch embedding layer, i.e., $f_{e}:\mathbb{R}^{N\times(C\cdot S)}\rightarrow\mathbb{R}^{S\times D}$ , where $D$ is the embedding dimension. The extracted features are denoted as $x_{e}$, and $x_{e}=f_{e}(I_{p})\in\mathbb{R}^{S\times D}$. When employing prompt tuning methods, prompts can be fed into model in various ways. For example, a prompt $P_{e}\in\mathbb{R}^{L_{p}\times D}$ is considered as an additional token concatenated to the embedding future $x_{e}$, where $L_{p}$ is the token length. In a remarkable work called Explicit Visual Prompting (EVP) [evp], prompts are first adjusted by a tunable adapter, then added to the extracted embedding feature $x_{e}$ through element-wise addition. This kind of prompt tuning typically yields better results because it learns explicit prompts. In this work, we adopt a approach similar to EVP.
@@ III-B Overview
As illustrated in Figure 2, Light-MLD employs a transformer-based backbone as encoder, which captures multi-stage features via several vision transformer blocks, then followed by a series of lightweight decoders to decode the corresponding landmarks for different tasks. Figure 3 gives a overview our proposed Adaptive Query Prompting (AQP). Prompts are selected from the prompt pool to instruct the model during training process.
@@ III-C Adaptive query prompting
In our proposed adaptive query prompting, there are three main components, namely the Prompt pool, the Prompt query mechanism and the Adaptor. It is worth mentioning that AQP only adds negligible additional parameters ($\sim$ 0.1%). Although AQP introduces extra prompts, its parameters are remarkably intricate, embodying a fundamentally different design principle from architecture-based methods: AQP designs a novel memory mechanism based on prompts, learning high-level instructions from model inputs to guide model outputs while maintaining the learned architecture unchanged. In contrast, most architecture-based methods aim to separate model parameters.
@@ III-C1 Prompt pool
where $N$ represents total number of prompts. $P_{j}\in\mathbb{R}^{L_{p}\times D}$ is a single prompt with the same embedding size $D$ as $x_{e}$ and token length $L_{p}$. $x_{e}$ is extracted embedding feature projected by the embedding layer. During the training phase, the prompt pool is randomly initialized.
@@ III-C2 Prompt query mechanism
Let $\gamma:\mathbb{R}^{D_{k}}\times\mathbb{R}^{D_{k}}\rightarrow\mathbb{R}$ represents a function that scores the match between the query and prompt key (in this work we adopt cosine distance). When presented with an input $x$, we use $q(x)$ to retrieve the top-M keys by straightforwardly solving the objective function:
@@ III-C3 Adaptor
$P^{i}_{f}=\operatorname{MLP_{up}}(\operatorname{GELU}(\operatorname{MLP^{i}_{tune}}(P))),$ (5)
where $MLP_{up}$ represents an up-projection layer shared across all Adaptors, aimed at aligning the dimensions of transformer feature. $GELU$ denotes the GELU activation function [bib53]. $MLP_{tune}$ is a linear layer responsible for generating distinct prompts in each Adaptor. $P^{i}_{f}$ is the output prompt that attaches to each transformer layer. This design allows for strong adaptability while generating diverse prompts. With the adaptor, we can generate suitable prompts for all tasks in a unified paradigm even when using different models or backbones.
@@ III-D2 Loss function
[…]
@@ IV-A1 Head
The head dataset is a widely-used public dataset for cephalometric landmark detection, which contains 400 X-ray images, provided in IEEE ISBI 2015 challenge [bib42]. It contains 19 landmarks for each X-ray image. For our experimentation, we designate the first 150 images for training and the remaining 250 images for testing purposes.
@@ IV-A2 Hand
The hand dataset comprises 909 X-ray images. We allocate the first 609 images for training and reserve the remaining 300 images for testing. Payer et al. [bib43] manually labeled a total of 37 landmarks.
@@ IV-A3 Chest
Models MRE Head SDR(%) MRE Hand SDR(%) MRE Chest SDR(%) (mm) 2mm 2.5mm 3mm 4mm (mm) 2mm 4mm 10mm (px) 3px 6px 9px Ibragimov et al. [bib47] 1.84 68.13 74.63 79.77 86.87 - - - - - - - - Štern et al. [bib52] - - - - - 0.80 92.20 98.45 99.83 - - - - Lindner et al. [bib48] 1.67 70.65 76.93 82.17 89.85 0.85 93.68 98.95 99.94 - - - - Urschler et al. [bib49] - 70.21 76.95 82.08 89.01 0.80 92.19 98.46 99.95 - - - - Payer et al. [bib50] - 73.33 78.76 83.24 89.75 0.66 94.99 99.27 99.99 - - - - U-Net [unet]+ 12.45 52.08 60.04 66.54 73.68 6.14 81.16 92.46 93.76 5.64 51.67 82.33 90.67 GU2Net [yolo]+ 1.56 77.79 84.65 89.41 94.93 0.84 95.40 99.35 99.75 5.57 57.33 82.67 89.33 AQP (Ours)+ 1.54 76.39 84.43 89.61 95.02 0.76 96.27 99.45 99.79 5.61 56.78 83.46 90.04
Figure 4: Subjective results of the head, hand and chest datasets. All images are randomly selected. The red points $\bullet$ are the landmarks predicted by our model while the green points $\bullet$ are the ground truth labels.
@@ IV-B Settings
All experiments are conducted on a single NVIDIA A100 GPU with 40G of memory. For model training, we adopt the AdamW optimizer with an initial learning rate of $\text{5e}^{-4}$. Cosine decay is employed to adjust the learning rate over time. We use the ViTPose-L as backbones for test, which are pre-trained on MS COCO Keypoint Detection Dataset [coco], corresponding to ViT-Large. The models are trained 25 epochs for all datasets. The images are resized to the shape of $256\times 192$ for all datasets.